OpenStudy (anonymous):

Will fan + give medal What is the 7th term of the geometric sequence where a1 = -5 and a5 = -1,280?

OpenStudy (anonymous):

geometric sequence formula is \(\sf \Large a_n=a_1 \times r^{n-1}\) first thing you have to do is to find the r or the ratio to find it, plug in your given values into the equation. \(\sf \Large a_n=a_1 \times r^{n-1}\\ -1280=-5 \times r^{5-1}\) simplify and solve for r, what will you get ?

OpenStudy (anonymous):

256?

OpenStudy (anonymous):

\(\sf \Large a_n=a_1 \times r^{n-1}\\ -1280=-5 \times r^{5-1}\\-1280\div -5=r^{4}\\256=r^4\\ r= 4\) you forget to evaluate \(\sf r^4\) :) now you have the value for r you can easily solve for \(\sf a_7\) using the formula

OpenStudy (anonymous):

Is it -20,480?

OpenStudy (anonymous):

\(\sf \Large a_n=a_1 \times r^{n-1}\\a_7=-5 \times (4)^{6}\\a_7=-5 \times 4096\\ a_7=-20480 \) looks right ^_^